OpenStudy (anonymous):

There are four hydrogen atoms in a molecule of methane (CH4). How many hydrogen atoms are there in two moles of methane? I know that the answer is 4.82 * 10^24 but I'm not completely sure how to get there

OpenStudy (aaronq):

so in two moles of methane (\(CH_4\)) there are 8 moles of hydrogen atoms, correct? Use: \(\sf moles=\dfrac{atoms}{N_A}\) where \(N_A\) is avogadro's number

OpenStudy (anonymous):

? mols=8/6.02x10^23=4.82x10^24

OpenStudy (aaronq):

you dont know the amount of atoms though, that's the unknown. 8 \( \sf \large moles\) of hydrogen atoms

OpenStudy (anonymous):

It says there are 4 atoms in a molecule

OpenStudy (aaronq):

yeah, in 1 molecule. later it says "in two moles of methane, how many hydrogen atoms?"

OpenStudy (anonymous):

So the problem is 2 moles=? atoms/6.02x10^23

OpenStudy (aaronq):

no. it's asking for the number of hydrogen atoms, not the number of molecules of methane. Think about this. you have two moles of methane \(CH_4\), but each molecule has 4 atoms of H. So there are 2 moles*4=8 moles of H atoms

OpenStudy (anonymous):

Each molecule of methane has 4 hydrogen atoms not each mole has 4 atoms

OpenStudy (aaronq):

lol did i say "each mole has 4 atoms"?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

I've really confused myself now. Obviously it isn't a good idea to go back and forth between stoichiometry and college applications

OpenStudy (aaronq):

I didn't say that. And no, you should probably focus on 1 thing at a time, specially if the concepts arent solid. I think your confusion is with what a mole is. A mole is just a number, like say a dozen - it's just a really large number. 1 dozen of eggs = 12 eggs 1 mole of eggs = 6.022*10^23 eggs 1 dozen methane molecules = 12 methane molecules each molecule has 4 atoms of hydrogen 1 dozen methane molecules have 4 dozens of hydrogen atoms it's the same with moles. 2 moles of methane have 8 moles of hydrogen (atoms)

OpenStudy (anonymous):

Okay. How does it get to 4.82*10^24? It gave us the correct answers but we have to come up with the calculations

OpenStudy (aaronq):

right. so we have established that there are 4 moles of H, right? now it's a plug and chug \(\sf 8~moles~of~H=\dfrac{\#~atoms}{N_A}\rightarrow \#~atoms=8~moles*N_A=4.8176*10^{24}\)

OpenStudy (aaronq):

i meant 4 moles per 1 mole of methane (thus 8 moles of H in total)

OpenStudy (anonymous):

Wouldn't the number of atoms be 8?

OpenStudy (aaronq):

why would it be?

OpenStudy (anonymous):

Because there are 4 atoms in each molecule and doesn't one molecule= one mole which mean that there are 8 atoms in two moles

OpenStudy (aaronq):

nope. 1 molecule is one molecule 1 mole of molecules = 6.022*10^23 molecules

OpenStudy (anonymous):

Okay